OpenStudy (anonymous):

if each interior angle of a regular polygon measures 144 degrees, how many sides does the polygon have?

OpenStudy (anonymous):

Please show steps on how to solve.

OpenStudy (anonymous):

obtuse angles but idont know about a polygon with obtuse angles

OpenStudy (jamesj):

Ok. What's the sum of angles of a triangle? 180 degrees. A triangle is a polygon with 3 sides. What's the sum of the angles of a square? A square is a polygon with 4 sides. What's the answer here?

OpenStudy (anonymous):

the answer is supposed to be 10 sides how do you solve that?

OpenStudy (anonymous):

decagon?

OpenStudy (jamesj):

We'll get there ... follow me in logic first

OpenStudy (anonymous):

ok

OpenStudy (jamesj):

What's the sum of angles of a square?

OpenStudy (anonymous):

180 (4-2) = 360

OpenStudy (anonymous):

360 ofcourse.as all angles are 90

OpenStudy (anonymous):

the polygon in question has 10 sides

OpenStudy (anonymous):

^ why?

OpenStudy (jamesj):

Right ... so you've got the formula for an n-sided polygon. It has a sum of angles \[ 180(n-2) \] and that's because you can make a polygon out of triangles. Stick two triangles together and you have a square, hence the sum of their angles is 180 + 180. Take a square add a triangle and you have a pentagon.

OpenStudy (anonymous):

i know how to do that, how do you solve when you dont know the number of sides?

OpenStudy (jamesj):

Now, if an n-sided polygon has a sum of angles 180(n-2), what's the size of the angles if each of them is the same? Well, there are n angles, hence one angle has the size of \[ \frac{180(n-2)}{n} \]

OpenStudy (jamesj):

For your problem, set that equal to 144 and solve for n.

OpenStudy (anonymous):

the general rule is Each Angle (of a Regular Polygon) = (n-2) × 180° / n so 144=(n-2)x180/n 144n=(n-2)x180 144n=180n-360 144n-180n=-360 36n=360 n=10

OpenStudy (anonymous):

so, James J, i do 180 (n-2) / n = 144 144-180+2= n/n?

OpenStudy (anonymous):

NOPE THATS NOT RIGHT ^

OpenStudy (anonymous):

i dont understnad..

OpenStudy (jamesj):

Safiah has written this out for you. But to start it off again \[ \frac{180(n-2)}{n} = 144 \] hence \[ 180(n-2) = 144n \] Can you do it now?

OpenStudy (anonymous):

OH WOW I GET IT!

OpenStudy (anonymous):

sooo 180( n-2) = 144n wiat no, what do i do now?

OpenStudy (jamesj):

...hence \[ 180n - 360 = 144n \] Now?

OpenStudy (anonymous):

expand the brackets hellobudy

OpenStudy (anonymous):

nope.. now what?

OpenStudy (jamesj):

Subtract 144n from both sides... \[ 180n - 144n - 360 = 0 \] i.e. \[ 36n - 360 = 0 \] Now can you finish it?

OpenStudy (anonymous):

-360 = -36 and then you divide 360 by 36 and you get ten

OpenStudy (jamesj):

right, \[ 36n = 360 \] hence \[ n = \frac{360}{36} = 10 \]

OpenStudy (anonymous):

k i think i get it..

OpenStudy (anonymous):

thank you!

OpenStudy (jamesj):

Do yourself a favor. Take a blank piece of paper. Write out the solution again. When you can do that without looking at anything --such as this web or another version of the solution--then you know you understand the solution.

OpenStudy (anonymous):

ok! can you give me another example so i can practice? use a different degrees?

OpenStudy (anonymous):

ok ill try with 100 degrees

OpenStudy (jamesj):

No, that won't work. One sec.

OpenStudy (jamesj):

try 108 degrees

OpenStudy (anonymous):

ALRIGHT :)

OpenStudy (anonymous):

5 sides?

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

:D

OpenStudy (jamesj):

last one, 162 degrees

OpenStudy (anonymous):

20 :)

OpenStudy (jamesj):

what about for 180 degrees?

OpenStudy (anonymous):

huh?

OpenStudy (jamesj):

That would correspond to a regular polygon with an infinite number of sides. What does that look like?

OpenStudy (anonymous):

idk waht your asking

OpenStudy (jamesj):

It would look like a circle.

OpenStudy (jamesj):

As you add more and more sides to a regular polygon, it looks more and more like a circle. In the limit, as the number of sides goes to infinity, it would become a circle. http://www.ck12.org/ck12/images?id=301055

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

g2g

OpenStudy (jamesj):

Here's an even better picture of this idea:

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Great job JamesJ - and way to hang in there hellobuddy!